OpenStudy (anonymous):

2/5+1/6 common denominator is 30

OpenStudy (anonymous):

correct LCD is 30

OpenStudy (anonymous):

17/30

OpenStudy (anonymous):

2*6=30 1*5=5 12/30 5/30=17/30 does it need to be reduced

OpenStudy (diyadiya):

17/30 does not have any common factor ,so it can't be reduced